# Global identification of dynamic panel models with interactive effects

### v1

@@ Abstract
This paper examines the problem of global identification in dynamic panel models with interactive effects, a fundamental issue in econometric theory. We focus on the setting where the number of cross-sectional units ($N$) is large, but the time dimension ($T$) remains fixed. While local identification based on the Jacobian matrix is well understood and relatively straightforward to establish, achieving global identification remains a significant challenge. Under a set of mild and easily satisfied conditions, we demonstrate that the parameters of the model are globally identified, ensuring that no two distinct parameter values generate the same probability distribution of the observed data. Our findings contribute to the broader literature on identification in panel data models and have important implications for empirical research that relies on interactive effects.
@@ 1 Introduction
$y_{it}=\alpha y_{it-1}+\delta_{t}+\lambda_{i}^{{}^{\prime}}f_{t}+\varepsilon_{it}$ (1)
@@
where $i=1,2,\dots,N$ and $t=1,2,\dots,T$. $y_{it}$ is the outcome variable, $\alpha$ is the autoregressive parameter, whereas $\delta_{t}$ captures time-specific fixed effects. $f_{t}$ is an $\bar{r}\times 1$ vector containing the value of the $\bar{r}$ latent factors at time $t$, and $\lambda_{i}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$. $\varepsilon_{it}$ is the idiosyncratic error term which we assume to be heteroskedastic but not autocorrelated i.e. $\mathbb{E}[\varepsilon_{i,t}^{2}]=\sigma_{t}^{2}\;\text{and}\;\mathbb{E}[\varepsilon_{i,t}\varepsilon_{i,s}]=0\;$ for $t\neq s$, and independent from the interactive effects. In this model, $y_{it}$ is the only observable variable.
@@
This framework provides a flexible approach to estimate unobservable trends from the data. The model allows the common shocks (modeled by $f_{t}$) to have a heterogeneous effect in individuals (modeled by $\lambda_{i}$) on the outcome variable $y_{it}$. A notable special case is the standard additive fixed-effects model, where $y_{it}=\delta_{t}+\mu_{i}+\varepsilon_{it}$, which corresponds to $\alpha=0$ and a single constant factor $f_{t}=1$. Thus, this model generalizes the additive framework by allowing for richer dynamics and heterogeneity in the factor structure. Parameter estimation in this model can be conducted via quasi maximum likelihood based methods as in Hayakawa et al. (2023) and Bai (2024). This paper studies the problem of global identification of the model parameters.
@@
Following Rothenberg (1971), we say that a parameter point $\theta^{0}$ in the parametric space $\Theta$ is globally identifiable if there is no other $\theta$ in $\Theta$ which is observationally equivalent. Establishing the identifiability of a model’s parameters is crucial, as it ensures that the estimation procedure yields meaningful and interpretable results.
@@
Despite the apparent simplicity of the model, demonstrating parameter identifiability is a nontrivial task. The seminal work of Anderson and Rubin (1956) established that in the absence of the autoregressive term ($\alpha=0$), the parameters of the model are globally identified up to an indeterminacy given by an $\bar{r}\times\bar{r}$ rotation matrix, provided that $T\geq 2\bar{r}+1$.
@@
To the best of our knowledge, no existing study has established the global identifiability of the parameters in the model described by (1). Bai (2024) demonstrates that the parameters of this model are locally identified when $\bar{r}$ does not exceed the modified Ledermann bound (Ledermann, 1937). In contrast, Hayakawa et al. (2023) argues that global identification is unattainable in dynamic panel data models with interactive effects.
@@
We demonstrate that under standard assumptions in the factor analysis literature, the parameters of the model are almost surely globally identifiable when $T\geq 2(\bar{r}+1)$. The global identification is achieved using the first two moments, which is equivalent to global identification for the quasi Gaussian maximum likelihood method. This result is of particular importance, as it establishes a one-to-one correspondence between the population moments and the model parameters, despite high-degree of nonlinearity of the problem.
@@
This paper contributes to the literature on global identification and panel data models. The question of global identification in parametric models has been the focus of extensive research (see, for example, Shapiro (1985), Komunjer and Ng (2011), Komunjer (2012), and Kocikecki and Kolasa (2023)), as well as trying to understand when and why identification fails (Forneron, 2024). A comprehensive review of this literature is provided in Lewbel (2019).
@@
Panel data models have been a core focus in econometrics for many years, with extensive study and application over the past few decades. Notable monographs and textbooks on panel data include works by Arellano (2003), Baltagi (2008), Hsiao (2022), and Wooldridge (2010), among others. More recently, there has been significant research on dynamic panel models with interactive effects, including studies by Moon and Weidner (2017), Bai (2024), and Hayakawa et al. (2023) to name a few. However, these studies have not addressed the global identifiability of the model parameters, a gap this paper aims to fill.
@@
The remainder of the paper is organized as follows: Section 2 formalizes the identification problem in our context. Section 3 outlines the assumptions required for our proof and presents some useful mathematical results that will be instrumental in our analysis. Section 4 derives intermediate results and provides the main proof of global identification under the stated assumptions. Finally, section 5 provides some concluding remarks with a discussion of potential extensions of the paper.
@@ 2.1 The Model in Matrix Notation
Before formally defining global identification, it is useful to express the model in matrix notation. Suppose we observe $y_{i}=(y_{i1},y_{i2},\dots y_{i,T})^{\prime}$. We begin by projecting the initial observation $y_{i1}$ onto $[1,\lambda_{i}]$, which gives
$y_{i1}=\delta_{1}^{*}+\lambda_{i}^{{}^{\prime}}f_{1}^{*}+\varepsilon_{i1}^{*}$
@@
where $y_{i}=[y_{i1},y_{i2},\dots,y_{iT}]^{\prime}$ is a $T\times 1$ vector of individual observations for $t=1,\dots,T$, $\delta=[\delta_{1},\dots\delta_{T}]^{\prime}$ is a $T\times 1$ vector, and $\mathbf{F}=[f_{1},f_{2},\dots,f_{T}]^{\prime}$ is a $T\times\bar{r}$ matrix containing the $\bar{r}$ factors, $\lambda_{i}=[\lambda_{i1},\lambda_{i2},\dots,\lambda_{i\bar{r}}]^{\prime}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$, and $\varepsilon_{i}=[\varepsilon_{i1},\varepsilon_{i2},\dots,\varepsilon_{i,T}]^{\prime}$ is a $T\times 1$ vector of idiosyncratic errors. Furthermore,
@@
a $T\times T$ matrix. Let $\mathbf{\Gamma}=\mathbf{B^{-1}}.$ Pre-multiplying by $\mathbf{B^{-1}}$ allows us to rewrite the model as
@@
In the original model the errors had a diagonal covariance matrix given by $\mathbf{D}=\operatorname*{diag}(d_{1},d_{2},\dots,d_{T}):=\operatorname*{diag}(\sigma_{1}^{2},\dots,\sigma_{T}^{2})$ (for notational simplicty, we write $\sigma_{t}^{2}$ as $d_{t}$). After transformation, the idiosyncratic error $\mathbf{\Gamma}\varepsilon_{i}$ has covariance matrix $\mathbf{\Gamma}\mathbf{D}\mathbf{\Gamma}^{\prime}$, which is no longer diagonal. Let
@@
denote the sample covariance matrix of $\lambda_{i}$, where $\bar{\lambda}=\frac{1}{N}\sum_{n=1}^{N}\lambda_{i}$ and let $\mathbb{E}[\mathbf{\Psi}_{N}]=\mathbf{\Psi}$. Similarly, define
@@
as the sample covariance matrix of $y_{i}$, where $\bar{y}=\frac{1}{N}\sum_{n=1}^{N}y_{i}$.
@@
$\mathbb{E}[\mathbf{S}_{N}]=\mathbf{\Sigma}(\theta)=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\Gamma}\mathbf{\Omega}\mathbf{\Gamma}^{\prime}$ (3)
where $\theta=\text{vec}(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ represents the model parameters. The matrix $\mathbf{\Sigma}$ is a $T\times T$ symmetric positive definite (PD) matrix, $\mathbf{\Psi}$ is an $\bar{r}\times\bar{r}$ symmetric PD matrix, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is a $T\times T$ symmetric positive semi-definite (PSD) matrix of rank $\bar{r}$, and $\mathbf{D}$ is a $T\times T$ diagonal PD matrix. $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}$ is a $T\times T$ symmetric PD matrix corresponding to the expected value of sample variance of $y_{i}$ in a basic model without an autoregressive term $\alpha=0$ (Anderson and Rubin, 1956).
@@ 2.2 Global identification of the model
This paper aims to establish that the parameters of the model are globally identifiable from the quasi log-likelihood perspective, which is equivalent to demonstrating global identification from the first two moments. Specifically, this means that the expectation and variance of $y_{i}$ uniquely determine the structural parameters of the model, as there exists a one-to-one mapping between them.
To formalize this argument, define $\delta^{\dagger}=\bm{\Gamma}\delta$. The quasi Gaussian log-likelihood function is given by
$-\frac{N}{2}\ln|\bm{\Sigma}(\theta)|-\frac{1}{2}\sum_{i=1}^{N}(y_{i}-\delta^{\dagger})^{\prime}\bm{\Sigma}(\theta)^{-1}(y_{i}-\delta^{\dagger})$
Concentrating out $\delta^{\dagger}$, which is estimated by $\bar{y}$, and dividing by $N$, we obtain [^1] In the MLE, the definition of $\mathbf{S}_{N}$ needs a slight modification, namely, $1/(N-1)$ is replaced by $1/N$.
$\frac{1}{N}\ell_{N}(\theta)=-\frac{1}{2}\ln|\bm{\Sigma}(\theta)|-\frac{1}{2}\mathrm{tr}[\bm{\Sigma}(\theta)^{-1}\mathbf{S}_{N}]$
[…]
Furthermore, it is known (Arnold, 1981) that this function has a unique maximizer at $\mathbf{\Sigma}(\theta)=\mathbf{\Sigma}(\theta^{0})$. While this result ensures that $\theta^{0}$ is a global maximizer of the objective function, it does not immediately establish uniqueness—hence, global identification—because it does not preclude the possibility that other parameter values $\widetilde{\theta}\neq\theta^{0}$ satisfy $\mathbf{\Sigma(\widetilde{\theta})}=\mathbf{\Sigma}(\theta^{0})$. In the following sections, we investigate whether any alternative parameter set $\widetilde{\theta}\neq\theta^{0}$ satisfies $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$. We demonstrate that, almost surely, no such alternative parametrization exists. Consequently, the quasi log-likelihood function is uniquely maximized at the true parameter value $\theta^{0}$, establishing global identification. This result, which has not been previously shown and was thought to be untrue (Hayakawa et al., 2023), constitutes a novel contribution.
@@
  • $\widetilde{\mathbf{\Gamma}}$ the counterpart to $\mathbf{\Gamma}$ under an alternative parameter $\widetilde{\alpha}$, given by
@@
We say that the parameters in the dynamic panel data model with interactive effects are globally identified at $\theta=\text{vec}(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ if there does not exist a diagonal and PD matrix $\mathbf{\widetilde{D}}$, a parameter $\widetilde{\alpha}$, and a symmetric PSD matrix $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ of rank $\bar{r}$, with at least one of them different from $\mathbf{D}$, $\alpha$, and $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$, respectively, such that
$\mathbf{\Sigma}=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\widetilde{\Gamma}}(\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}})\mathbf{\widetilde{\Gamma}}^{\prime}$
Equivalently, we want to show
@@
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$, $\mathbf{\widetilde{\Gamma}}=\mathbf{\Gamma}$ (or equivalently $\widetilde{\alpha}=\alpha$), and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}$.

### v2

@@ Abstract
We investigate the problem of global identification in dynamic panel models with interactive effects, under the large-$N$, fixed-$T$ setting. While local identification, typically established via the Jacobian matrix, is well understood, global identification has remained a more elusive and challenging issue. It is commonly believed to be unachievable in this context. However, we demonstrate that the model is, in fact, globally identified. Our analysis also covers models with additive fixed effects, including cases with unit roots, where previous studies have reported non-identification based on certain estimators. We show that, even in these settings, global identification holds.
@@ 1 Introduction
$y_{it}=\alpha y_{it-1}+\delta_{t}+\lambda_{i}^{\prime}f_{t}+\varepsilon_{it}$ (1)
@@
where $i=1,2,\dots,N$ and $t=1,2,\dots,T$. $y_{it}$ is the outcome variable, $\alpha$ is the autoregressive parameter, whereas $\delta_{t}$ captures time-specific fixed effects. $f_{t}$ is an $\bar{r}\times 1$ vector containing the value of the $\bar{r}$ latent factors at time $t$, and $\lambda_{i}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$. $\varepsilon_{it}$ is the idiosyncratic error term, which we assume to be i.i.d. across $i$, heteroskedastic across $t$ but serially uncorrelated, i.e. $\mathbb{E}[\varepsilon_{it}^{2}]=\sigma_{t}^{2}\;\text{and}\;\mathbb{E}[\varepsilon_{it}\varepsilon_{is}]=0\;$ for $t\neq s$, and independent from the interactive effects. In this model, $y_{it}$ is the only observable variable.
@@
This framework provides a flexible way to model unobservable common trends and heterogeneous responses. The model allows the common shocks (modeled by $f_{t}$) to have a heterogeneous effect in individuals (modeled by $\lambda_{i}$) on the outcome variable $y_{it}$. A notable special case is the standard additive fixed-effects model, where $y_{it}=\delta_{t}+\gamma_{i}+\varepsilon_{it}$, which corresponds to $\alpha=0$ and a single constant factor $f_{t}=1$. Thus, this model generalizes the additive framework by allowing for richer dynamics and heterogeneity in the factor structure. Parameter estimation in this model can be conducted via quasi maximum likelihood based methods as in Hayakawa et al., (2023) and Bai, (2024).
This paper studies the problem of global identification of the model parameters when the number of cross-sectional units $(N)$ is large, but the time dimension $(T)$ is fixed. This case is particularly relevant in applied economics because many panel datasets, such as those involving countries, firms, or households, span a large number of cross-sectional units but only cover a limited time horizon. For instance, macroeconomic panels often include many countries but only a few decades of annual data, making the short-$T$, large-NN framework empirically realistic.
@@
Following Rothenberg, (1971), we say that a parameter point $\theta^{0}$ in the parametric space $\Theta$ is globally identifiable if there is no other $\widetilde{\theta}\in\Theta$ which is observationally equivalent. Establishing the identifiability of a model’s parameters is crucial, as it ensures that the estimation procedure yields meaningful and interpretable results.
@@
Despite the apparent simplicity of the model, demonstrating parameter identifiability is a nontrivial task. The seminal work of Anderson and Rubin, (1956) established that in the absence of the autoregressive term ($\alpha=0$), the parameters of the model are globally identified up to an indeterminacy given by an $\bar{r}\times\bar{r}$ rotation matrix, provided that $T\geq 2\bar{r}+1$.
@@
To the best of our knowledge, no existing study has established the global identifiability of the parameters in the model described by (1). Bai, (2024) demonstrates that the parameters of this model are locally identified when $\bar{r}$ does not exceed the modified Ledermann bound (Ledermann,, 1937). In contrast, Hayakawa et al., (2023) argues that global identification is unattainable in dynamic panel data models with interactive effects.
@@
We demonstrate that under standard assumptions in the factor analysis literature, the parameters of the model are almost surely globally identifiable when $T\geq 2(\bar{r}+1)$. The global identification is achieved using the first two moments. Since the quasi-Gaussian likelihood depends only on the first two moments, this equivalence implies global identification for quasi-Gaussian maximum likelihood estimation. This result is of particular importance, as it establishes a one-to-one correspondence between the population moments and the model parameters, despite high degree of nonlinearity of the problem.
@@
This paper contributes to the literature on global identification and panel data models. The question of global identification in parametric models has been the focus of extensive research (see, for example, Shapiro, (1985), Komunjer and Ng, (2011), Komunjer, (2012), and Kociecki and Kołasa, (2023)), while other authors have focused on trying to understand when and why identification fails (Forneron,, 2024). A comprehensive review of this literature is provided in Lewbel, (2019).
@@
Panel data models have been a core focus in econometrics for many years, with extensive study and application over the past few decades. Notable monographs and textbooks on panel data include works by Arellano, (2003), Baltagi, (2008), Hsiao, (2022), and Wooldridge, (2010), among others. Moreover, the study of panel data models with interactive fixed effects [^1] Sometimes referred to in the literature as multiple time-varying individual effects (Ahn et al.,, 2013) or multifactor error structure (Pesaran,, 2006). has gained notable attention during the recent years. While there is an extensive literature in this area, some of the most noteworthy papers include Holtz-Eakin et al., (1988), Ahn et al., (2001), Pesaran, (2006), Bai, (2009), and Ahn et al., (2013). More recently, there has been significant research on dynamic panel models with interactive effects, including studies by Moon and Weidner, (2017), Hayakawa et al., (2023), and Bai, (2024), to name a few. However, these studies have not addressed the global identifiability of the model parameters for the short-$T$ case, a gap this paper aims to fill.
We also argue that the techniques developed in this paper can be used to prove identification in a broad class of dynamic models featuring time, individual and interactive effects. For example, we show that in a dynamic panel with individual effects (Arellano and Bond,, 1991), the autoregressive coefficient can be globally identified even when $\alpha=1$, a case previously thought to be non-identifiable (see Blundell and Bond, 1998; Sentana, 2024).
@@
The remainder of the paper is organized as follows: Section 2 formalizes the identification problem in our context. Section 3 outlines the assumptions required for our proof and presents some useful mathematical results that will be instrumental in our analysis. Section 4 derives intermediate results and provides the main proof of global identification under the stated assumptions. Section 5 discusses identification in the presence of individual fixed effects. Finally, Section 6 provides some concluding remarks with a discussion of potential extensions of the paper.
@@ 2.1 The Model in Matrix Notation
Before formally defining global identification, it is useful to express the model in matrix notation. Suppose we observe $y_{i}=(y_{i1},y_{i2},\dots,y_{iT})^{\prime}$. We begin by projecting the initial observation $y_{i1}$ onto $1,\lambda_{i}]$, which gives
$y_{i1}=\delta_{1}^{*}+\lambda_{i}^{\prime}f_{1}^{*}+\varepsilon_{i1}^{*}$
@@
where $y_{i}=[y_{i1},y_{i2},\dots,y_{iT}]^{\prime}$ is a $T\times 1$ vector of individual observations for $t=1,\dots,T$, $\delta=[\delta_{1},\dots,\delta_{T}]^{\prime}$ is a $T\times 1$ vector, and $\mathbf{F}=[f_{1},f_{2},\dots,f_{T}]^{\prime}$ is a $T\times\bar{r}$ matrix containing the $\bar{r}$ factors, $\lambda_{i}=[\lambda_{i1},\lambda_{i2},\dots,\lambda_{i\bar{r}}]^{\prime}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$, and $\varepsilon_{i}=[\varepsilon_{i1},\varepsilon_{i2},\dots,\varepsilon_{iT}]^{\prime}$ is a $T\times 1$ vector of idiosyncratic errors. Furthermore,
@@
is a nonsingular lower-triangular $T\times T$ matrix, hence invertible. Let $\mathbf{\Gamma}=\mathbf{B}^{-1}.$ Pre-multiplying by $\mathbf{B}^{-1}$ allows us to rewrite the model as
@@
In the original model, the errors had a diagonal covariance matrix given by $\mathbf{D}=\operatorname*{diag}(d_{1},d_{2},\dots,d_{T}):=\operatorname*{diag}(\sigma_{1}^{2},\dots,\sigma_{T}^{2})$ (for notational simplicity, we write $\sigma_{t}^{2}$ as $d_{t}$). After transformation, the idiosyncratic error $\mathbf{\Gamma}\varepsilon_{i}$ has covariance matrix $\mathbf{\Gamma}\mathbf{D}\mathbf{\Gamma}^{\prime}$, which is no longer diagonal. Let
@@
denote the sample covariance matrix of $\lambda_{i}$, where $\bar{\lambda}=\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}$ and let $\mathbb{E}[\mathbf{\Psi}_{N}]=\mathbf{\Psi}$. Similarly, define
@@
as the sample covariance matrix of $y_{i}$, where $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$.
@@
$\mathbb{E}[\mathbf{S}_{N}]=\mathbf{\Sigma}(\theta^{0})=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\Gamma}\mathbf{\Omega}\mathbf{\Gamma}^{\prime}$ (3)
where $\theta^{0}=(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ is the true vector of structural parameters. The matrix $\mathbf{\Sigma}$ is a $T\times T$ symmetric positive definite (PD) matrix, $\mathbf{\Psi}$ is an $\bar{r}\times\bar{r}$ symmetric PD matrix, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is a $T\times T$ symmetric positive semi-definite (PSD) matrix of rank $\bar{r}$, and $\mathbf{D}$ is a $T\times T$ diagonal PD matrix. $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}$ is a $T\times T$ symmetric PD matrix corresponding to the expected value of the sample covariance matrix of $y_{i}$ in a basic model without an autoregressive term $\alpha=0$ (Anderson and Rubin,, 1956).
@@ 2.2 Global identification of the model
This paper aims to establish that the parameters of the model are globally identifiable from the Gaussian quasi log-likelihood perspective, which is equivalent to demonstrating global identification from the first two moments. Specifically, this means that the expectation and variance of $y_{i}$ uniquely determine the structural parameters of the model, as there exists a one-to-one mapping between them.
To formalize this argument, define $\delta^{\dagger}=\mathbf{\Gamma}\delta$. The Gaussian quasi log-likelihood function is given by
$\frac{N}{2}\ln|\mathbf{\Sigma}(\theta)|-\frac{1}{2}\sum_{i=1}^{N}(y_{i}-\delta^{\dagger})^{\prime}\mathbf{\Sigma}(\theta)^{-1}(y_{i}-\delta^{\dagger})$
Concentrating out $\delta^{\dagger}$, which is estimated by $\bar{y}$, and dividing by $N$, we obtain [^2] In the MLE, the definition of $\mathbf{S}_{N}$ needs a slight modification, namely, 1/(N1)1/(N-1) is replaced by $1/N$.
$\frac{1}{N}\ell_{N}(\theta)=-\frac{1}{2}\ln|\mathbf{\Sigma}(\theta)|-\frac{1}{2}\mathrm{tr}[\mathbf{\Sigma}(\theta)^{-1}\mathbf{S}_{N}]$
For a fixed $T$ and as $N\to\infty$, by the law of large numbers, $\mathbf{S}_{N}\overset{p}{\to}\mathbf{\Sigma}(\theta^{0})$, we have
$\frac{2}{N}\ell_{N}(\theta)\overset{p}{\to}-\ln|\mathbf{\Sigma}(\theta)|-\mathrm{tr}\left[\mathbf{\Sigma}(\theta^{0})\mathbf{\Sigma}(\theta)^{-1}\right]$ (4)
Furthermore, it is known (Arnold,, 1981) that this function has a unique maximizer at $\mathbf{\Sigma}(\theta)=\mathbf{\Sigma}(\theta^{0})$. While this result ensures that $\theta^{0}$ is a global maximizer of the objective function, it does not immediately establish uniqueness—hence, global identification—because it does not preclude the possibility that other parameter values $\widetilde{\theta}\neq\theta^{0}$ satisfy $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$. In the following sections, we investigate whether any alternative parameter set $\widetilde{\theta}\neq\theta^{0}$ satisfies $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$. We demonstrate that, almost surely, no such alternative parametrization exists. Consequently, the quasi log-likelihood function is uniquely maximized at the true parameter value $\theta^{0}$, establishing global identification. This result, which has not been previously shown and was thought to be untrue (Hayakawa et al.,, 2023), constitutes a novel contribution.
@@
  • $\widetilde{\mathbf{\Gamma}}$ the analogue of $\mathbf{\Gamma}$ under an alternative parameter $\widetilde{\alpha}$, given by
@@
The model is globally identified at $\theta^{0}$ if whenever $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$, it must hold that $\widetilde{\alpha}=\alpha$, $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}\widetilde{\Psi}\widetilde{F}}’=\mathbf{F\Psi F’}$. Formally:
@@
If and only if $\widetilde{\alpha}=\alpha$, $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}\widetilde{\Psi}\widetilde{F}}’=\mathbf{F\Psi F’}$ (or simply $\widetilde{\theta}=\theta^{0}$).